OpenStudy (anonymous):

State the excluded values of a and b in the expression (3a^2b)/(2a^3b) 1) a ≠ 0 2) a≠2, b≠3 3) a≠0, b≠0 4) a≠3, b≠2

OpenStudy (anonymous):

\[\left(3 a^2 b\right)/\left(2 a^3 b\right)=\frac{3}{2 a} \]\[a\neq 0 \]

OpenStudy (anonymous):

Thank you for the medal.

OpenStudy (anonymous):

Thanks!